OpenStudy (anonymous):

MEDAL AND FAN :D Part A: The area of a square is (9x2 − 12x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (25x2 − 16y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points)

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

9x^2 - 12 x + 4 is a perfect square can you tell me what are the square roots of 9x^2 and 4?

OpenStudy (anonymous):

Idk, 4 square root is 2 right?

OpenStudy (welshfella):

yes what about 9x^2?

OpenStudy (welshfella):

u need square root of 9 and square root of x^2

OpenStudy (anonymous):

3, idk how to get the square root of x^2

OpenStudy (welshfella):

its easy x^2 = x * x so by definition square root of x^2 = x so square root of 9x^2 = 3x

OpenStudy (welshfella):

we need -2 for square root of 4 because we have -12x in the middle of the expression square root of 9x^2 - 12 x + 4 = (3x - 2) x^2 - 12 x + 4 = (3x - 2)^2 = area of the square each side of square has length 3x - 2

OpenStudy (anonymous):

Makes sense

OpenStudy (welshfella):

thats 9x^2 on last line not x^2

OpenStudy (anonymous):

So what would that get for part A?

OpenStudy (welshfella):

Part B is a rectangle with different length and width

OpenStudy (welshfella):

- we finished part A

OpenStudy (anonymous):

Wait, what was the answer for part a? x^2 - 12 x + 4 = (3x - 2)^2 ?

OpenStudy (welshfella):

yea those are the factors and side length = 3x - 2

OpenStudy (anonymous):

Sweet, now what would part b be? how would we find it?

OpenStudy (welshfella):

25x^2 - 16y^2 is the Difference of 2 squares this factors to 2 binomials which are the same except one has + and one has - sign first what are the sqqaure roots of 25x^2 and 16y^2?

OpenStudy (anonymous):

5x^2, 4x^2?

OpenStudy (welshfella):

first square root of 25 = 5 and sqrt of x^2 = x giving 5x so can you figure out square root of 16y^2 in a similar way?

OpenStudy (welshfella):

nearly right NOT 4x^2

OpenStudy (anonymous):

2x^2?

OpenStudy (welshfella):

you have the number paret right - 5 and 4 but what are sqrts of x^2 and y^2?

OpenStudy (anonymous):

x*x

OpenStudy (welshfella):

no x^2 = x*x - i wan the square root of x^2

OpenStudy (welshfella):

check back to earlier posts

OpenStudy (anonymous):

x?

OpenStudy (welshfella):

yes the square root of x^2 = x because x*x = x^2 similarly sqrt y^2 = y

OpenStudy (welshfella):

so sqrt 25x^2 = 5x and sqrt 16y^2 = ?

OpenStudy (anonymous):

2x?

OpenStudy (welshfella):

why 2x - there's no x in 16y^2 and sqrt 16 is not 2 - its 4

OpenStudy (anonymous):

so it's just 4

OpenStudy (welshfella):

4 is aqrt 16 yes but i need sqrt of 16y^2

OpenStudy (anonymous):

4x?

OpenStudy (welshfella):

- u r driving me crazy!! - how can it be x when theres no x in 16y^2 - what letter is in 16y^2?

OpenStudy (anonymous):

Square root of ^2 is x? waaait im wrong

OpenStudy (anonymous):

there is none, so why wouldn't it just be 4??

OpenStudy (anonymous):

4y!!!!!!!!??

OpenStudy (welshfella):

yes - of course 4y

OpenStudy (welshfella):

so going back 25x^2 - 16y^2 = ( 5x + 4y)(5x - 4y) note there's + in one bracket and - in the other

OpenStudy (welshfella):

those are the factors and the dimensions of the rectangle one dimension is 5x+4y and the other ix 5x-4y.

OpenStudy (anonymous):

So that's the answer, 5x+4y for each side?

OpenStudy (welshfella):

no 5x+4y for one side and 5x-4y for the other

OpenStudy (welshfella):

5x+4y will be the length and 5x-4y will be the width

OpenStudy (anonymous):

Thanks!!

OpenStudy (welshfella):

yw